OpenStudy (anonymous):

What are the coordinates of the center of the hyperbola?(x-3)^2/81 - y^2/144= 1

OpenStudy (zpupster):

equation of hyperbola \[\frac{ (x-h)^2 }{ a^2}-\frac{ (y-k)^2 }{ b^2 }=1\] center (h,k) watch your signs what did you get?

OpenStudy (anonymous):

(0,- 3)

OpenStudy (anonymous):

@zpupster

OpenStudy (loser66):

nope, it's (3,0)